A company produces three products P, Q and R using raw materials A, B and C. One unit of P requires 1, 2 and 3 units of A, B and C respectively. One unit of Q requires 2, 3 and 2 units of A, B and C respectively. One unit of R requires 1, 2 and 2 units of A, B and C respectively. The number of units available for raw material A, B and C are 8, 14 and 13 units respectively. Using the matrix method, determine the number of units of each product to produce in order to utilize completely the available resources.
1
Expert's answer
2020-11-17T17:24:12-0500

"\def\arraystretch{1.5}
 \begin{array}{c|c:c:c}
 & A(8) & B(14) & C(13)\\ \hline
 P & 1& 2& 3\\
 \hdashline
Q& 2 & 3& 2\\
 \hdashline
 R & 1& 2 & 2
\end{array}"

"A=\begin{pmatrix}
 1& 2 &1 \\ 
 2& 3 & 2\\ 
 3& 2 & 2
\end{pmatrix},"

"B=\begin{pmatrix}
 8 \\ 
 14\\ 
 13
\end{pmatrix},"

"X=\begin{pmatrix}
 P \\ 
 Q\\ 
 R
\end{pmatrix}."


"|A|=\begin{vmatrix}
 1& 2 &1 \\ 
 2& 3 & 2\\ 
 3& 2 & 2
\end{vmatrix}=1\cdot(6-4)-2\cdot(4-6)+1\cdot(4-9)=1\neq 0."


"a_{11}=\begin{vmatrix}
 3&2 \\ 
 2 & 2
\end{vmatrix}=2," "a_{12}=-\begin{vmatrix}
 2&2 \\ 
 3 & 2
\end{vmatrix}=2," "a_{13}=\begin{vmatrix}
 2&3 \\ 
 3 & 2
\end{vmatrix}=-5,"

"a_{21}=-\begin{vmatrix}
 2&1 \\ 
 2 & 2
\end{vmatrix}=-2," "a_{22}=\begin{vmatrix}
 1&1 \\ 
 3 & 2
\end{vmatrix}=-1," "a_{23}=-\begin{vmatrix}
 1&2 \\ 
 3 & 2
\end{vmatrix}=4,"

"a_{31}=\begin{vmatrix}
 2&1 \\ 
 3 & 2
\end{vmatrix}=1," "a_{32}=-\begin{vmatrix}
 1&1 \\ 
 2 & 2
\end{vmatrix}=0," "a_{33}=\begin{vmatrix}
 1&2 \\ 
 2 & 3
\end{vmatrix}=-1."


"adjA={\begin{pmatrix}
2 & 2 &-5 \\ 
-2 & -1 &4 \\ 
 1& 0& -1
\end{pmatrix}}^{T}={\begin{pmatrix}
 2& -2 &1 \\ 
 2& -1& 0\\ 
 -5&4 & -1
\end{pmatrix}}."


"A^{-1}=\frac{adjA}{|A|}=\begin{pmatrix}
 2& -2 &1 \\ 
 2& -1& 0\\ 
 -5&4 & -1
\end{pmatrix}."


"X=A^{-1}B."


"\begin{pmatrix}
 P \\ 
 Q\\ 
 R
\end{pmatrix}=\begin{pmatrix}
 2& -2 &1 \\ 
 2& -1& 0\\ 
 -5&4 & -1
\end{pmatrix}\cdot \begin{pmatrix}
 8 \\ 
 14\\ 
 13
\end{pmatrix}=\begin{pmatrix}
 16-28+13 \\ 
 16-14\\ 
 -40+56-13
\end{pmatrix}=\begin{pmatrix}
 1 \\ 
 2\\ 
 3
\end{pmatrix}."

"P=1, Q=2, R=3."